OpenStudy (anonymous):

Graph the function d=√1.5h which gives the approximate distance d in miles to the horizon, where h is the height of the viewer’s eyes above the ground in feet. Approximately how high above the ground must a person be if the horizon is 17.2 miles away?

OpenStudy (anonymous):

Please help. Don't understand at all.

OpenStudy (anonymous):

Ok, replace d with 17.2 to see what the value of x would be

OpenStudy (anonymous):

Sorry, not x, h

OpenStudy (anonymous):

To solve a balanced equation, we must get the variable by itself. To do that, we'll do the opposite operation. In this case, we see a \(\sqrt{1.5h}\). To get rid of the square root, we'll do the opposite operation, squaring. Of course, to keep the equation balanced, we will have to do the same thing to both sides.

OpenStudy (anonymous):

17.2 = sqrt(1.5*h) 17.2^2 = 1.5h h = (17.2^2)/1.5

OpenStudy (anonymous):

I'm still not getting it. The answer choices are a.3ft (Starts at 2 and increases) b.197ft c.3ft (Starts at 0 and icreases) 4.41ft

OpenStudy (anonymous):

@Chlorophyll Can you help?

OpenStudy (anonymous):

Did you follow others helpers' instructions?

OpenStudy (anonymous):

I have no idea what I'm doing. At all o.e

OpenStudy (anonymous):

The question seems wordy, but all it says distance d = √ (1.5h) with given d = 17.2 miles Therefore, all you do is to solve it, and the other helpers already set everything out for you!

OpenStudy (anonymous):

It's 197?

OpenStudy (anonymous):

Be careful about the units distance d in miles, but height h in feet!

OpenStudy (anonymous):

197 feet I meant. o;